# Wichitrnithed, August 28, 2025 GPU-acceleration of the Discontinuous Galerkin Shallow Water Equations Solver (DG-SWEM) using CUDA and OpenACC

### v1

@@ Abstract
This paper presents a porting of DG-SWEM, a discontinuous Galerkin solver for coastal ocean circulation, and in particular storm surge, to GPU using two separate approaches: CUDA Fortran and OpenACC. Time-explicit discontinuous Galerkin methods have been shown to exhibit a large amount of data parallelism due to the loose coupling between elements, and thus are naturally mapped to the GPU architecture. For each porting approach, we discuss the code design, ease of programming, and performance when running on realistic use cases. Specifically for the OpenACC version, we also aim to preserve maintainability within the same codebase through using Unified Memory. We test the codes on NVIDIA’s Grace Hopper chip and compare the GPU performance on each node to the MPI version on a single CPU node (144 cores).
@@ 1 Introduction
A primary disadvantage of the DG method which has confined DG-SWEM to mainly research and not operational forecasting like ADCIRC is the increased computational cost associated with the computations required from the higher number of global degrees of freedom compared to the continuous formulation. At the same time, the localized nature of computing the solution for each element independently generates a high degree of data parallelism, and several GPU implementations of the DG method have been using different approaches. Early implementations using NVIDIA’s CUDA include [Klöckner et al. (2009), Siebenborn et al. (2012), Fuhry (2013), Brodtkorb et al. (2012)]. More recent implementations using different application programming interfaces (API) can be seen in [Abdi et al. (2019), Kirby and Mavriplis (2020)] which use OCCA, and [Xia et al. (2014), Dai et al. (2022)] which use OpenACC. This paper presents and compares the porting of DG-SWEM to GPU using two separate approaches: CUDA and OpenACC. The aim of the latter approach is to explore a directive-based programming model which requires less integration and maintenance efforts, as well as to take advantage of the new NVIDIA Grace-Hopper architecture [Singer (2024)] which provides a more efficient CPU-GPU data transfer. Lastly, we mention the recent open-source FEniCSx SWE solverby Dawson et al. in [Dawson et al. (2024)] and GPU accelerated by Pachev in [Pachev (2024)], where implicit schemes are considered.
In the following, we introduce, discuss, and validate the proposed GPU acceleration. In section 2, we introduce the governing model, i.e., the SWE and our DG discretization. Next, in Sections 3 and 4, we discuss the acceleration approach and the hardware setup we utilize. In 6, we present numerical studies of our methodology including benchmark tests and large scale hurricanes that include compound flood effects. Finally, in Sections 7 and 8, we discuss our results, draw conclusions, and discuss potential future research directions.
@@ 3 GPU Programming approach
As an alternative to CPU parallelism, graphics processing units (GPUs) can be viewed as a multithreaded vector processor [Hennessy and Patterson (2018)], where each unit executes the same instructions on multiple data (SIMD), but individual lanes can take different paths. To take advantage of a large number of such processors, the algorithm must exhibit a high degree of data parallelism, where multiple elements can be processed independently at the same time. This is particularly applicable to explicit DG methods due to the inherently local computation of the elements.
In the past, as GPUs have originally been designed for graphics programming, the available application programming interface (API) such as DirectX and OpenGL imposed limitations of using GPUs for general-purpose computing (GPGPU) [Kirk and Hwu (2010)]. In 2007, NVIDIA developed CUDA [NVIDIA (2024)] which provides a more generic API to GPU instructions. This significantly extended the capabilities of NVIDIA GPUs to scientific and engineering codes: the programmer gained control of individual GPU threads, synchronization, and shared memory. Using the NVIDIA HPC compiler stack (a successor of PGI compilers), CUDA is supported in C, C++, and Fortran [NVIDIA (2020)]; CUDA code can be written in the same source files and compiled with appropriate flags. While the low-level nature of CUDA allows more opportunities for control and optimization by the programmer, it typically results in having to maintain a separate code base from the CPU code.
Another approach to GPU programming is to use OpenACC [OpenACC-Standard.org (2021)] and OpenMP [Openmp.org (2021)]
to write directives instead of manually managing individual GPU threads. This approach allows the programmer to insert high-level statements that generate GPU instructions and appear as comments in the source code (similar to OpenMP loop constructs, for instance). This can sometimes result in a single code base — only the compilation flags such as -acc need to be turned on to enable the GPU code generation. Of course, the programmer is still required to sufficiently expose the parallelism in the code, which often results in modifications of the CPU code as well.
In this work, we port DG-SWEM using each approach and compare their implementations and performance. For the directive-based approach, we choose to use OpenACC as it currently has better support on the NVHPC compilers.
@@ 4 Experimental setup
The original CPU code was compiled using the Intel oneAPI compiler stack. This work uses the NVIDIA SDK compiler stack (nvfortran, nvc) for both CPU and GPU versions, with the latter also using CUDA Fortran. Compilation flags for the CPU and GPU codes are listed in Table 3 . On Vista, the NVHPC version is 24.7 while on the local setup it is 23.11.
@@
| CUDA | nvfortran -r4 -O3 -cuda -Mextend -Mpreprocess - |
| OpenACC | nvfortran -r4 -O3 -cuda -acc -Mextend -Mpreprocess -gpu=unified |
@@ 5 Implementation
This section details our implementation for both CUDA Fortran and OpenACC for different aspects of the code. In the Kernels section, a general parallelization is outlined, then details specific to either approaches are explained.
5.1 Data management
A significant part of porting code to GPUs is managing data movement. As a CPU (host) and a GPU (device) have separate physical memories, data must be transferred back and forth to perform the respective calculations. Even if some arrays are only used in device calculations, they still need to be initialized or read from files through the host, and we would have to declare device copies of those arrays. For example, an existing array in the code
@@ 5.1 Data management
To avoid having to do this for each array in the code, we use CUDA’s unified and constant memory types, explained below.
5.1.1 Unified memory
Unified Virtual Addressing (UVA) was introduced in CUDA 6.0. This allows GPU kernels to access data on the host without needing to explicitly copy the data to device. Data migration is automatically performed once a page fault is triggered on the device. One way to achieve this is through the managed attribute. This way, we can reuse the existing arrays in the code and only modify their declarations, e.g.
real, allocatable, managed :: A(:,:)
! Use A on the device
istat = cudaDeviceSynchronize()
! Use A on the host
For newer systems that support Heterogenous Memory Management (HMM), the declaration can be completely eliminated using the compilation flag -gpu=unified. This significantly simplifies programming efforts and enhances maintainability. In both approaches, a synchronization is needed for correctness before the CPU can access data, since CUDA subroutines immediately return control to the host.
In this work, the CUDA Fortran approach uses the managed declarations to make explicit to the programmer which arrays will be used on the GPU, while the OpenACC code uses -gpu=unified to avoid any data declarations.
5.1.2 Constant memory
The standard way of passing scalar variables to CUDA kernels is by value, i.e. through the Fortran value attribute in the dummy argument declaration. When there are a lot of variables, such as constants, this increases the programmer’s efforts to pass them correctly. To avoid this, we use CUDA’s constant attribute for constant variables. They are declared and initialized in the host global module, and reside as read-only variables on the GPU constant memory partition which a size limit of 64 KB. Additionally, the constant memory has faster access time than global memory when all threads access the same element — a common pattern for many fixed-size arrays like quadrature weights. In Table 4, we summarize the important arrays whose attributes are modified, some of which are referred to in the kernel outline in the next section.
| Array | Meaning | Attribute |
| ZE[k,e,irk] | mode of elevation solution at element on RK stage irk | managed |
| QX[k,e,irk] | mode of x-velocity solution at element on RK stage irk | managed |
| QY[k,e,irk] | mode of y-velocity solution at element on RK stage irk | managed |
| RHS_ZE[k,e,irk] | Contribution to the mode of at element on RK stage irk | managed |
| RHS_QX[k,e,irk] | Contribution to the mode of at element on RK stage irk | managed |
| RHS_QY[k,e,irk] | Contribution to the mode of at element on RK stage irk | managed |
| PHI_AREA[k,i] | basis evaluated at area quadrature point i | constant |
| PHI_EDGE[k,i] | basis evaluated at edge quadrature point i | constant |
| XFAC[k,i,e] | Weight for the basis of element in x-direction at area quadrature point i | managed |
| YFAC[k,i,e] | Weight for the basis of element in x-direction at area quadrature point i | managed |
| SRFAC[k,i,e] | Weight for the basis of the source term on element at area quadrature point i | managed |
| EDGEFAC[k,i,l] | Weight for the basis of the source term on element at area quadrature point i | constant |
| EDGE_ze[k,l,e] | Numerical flux contribution to the basis of at local edge of element | managed |
| EDGE_qx[k,l,e] | Numerical flux contribution to the basis of at local edge of element | managed |
| EDGE_qy[k,l,e] | Numerical flux contribution to the basis of at local edge of element | managed |
@@ 5.2 GPU kernels
This section outlines the main subroutines and how they map to GPU parallelism. For the integration kernels, we follow the one-thread-per-element and one-thread-per-edge approach as in [Fuhry (2013)].
@@ 5.2.1 Area integration
The volume integration routine (Algorithm 1) is naturally data-parallel, as computations are done locally to each element . At a given Runge-Kutta stage irk, it computes the terms
from (3) and accumulates them into the RHS vectors. Each element is assigned to a thread identified by its global ID (in the CUDA version, this is the variable
(blockidx%x-1)*blockdim%x + threadidx%x)
which then computes the RHS contributions. This is done by looping through the quadrature points; for each quadrature point we loop through the degrees of freedom for each state variable and compute the flux terms. Recall from Eq 2 that these terms (explicitly written in terms of the state variables) are
@@
In the OpenACC version (Listing 2), the thread-level loop is enabled through the gang-vector parallelism over the number of elements (NE), and the inner loops over quadrature points and degrees of freedom are enforced to be serial. This needs to be explicitly inserted, as by default the inner loops will also be parallelized, and in this case we know the trip counts is too short to be profitable. Importantly, the async clause is added to eliminate kernel launch overhead, and the default(present) the overhead of checking if data is already on the device. This same structure is applied for the edge integration kernel. We have observed that using these clauses speeds up the program by almost 1.5.
@@ 5.2.2 Edge integration
In Algorithm 2, each CUDA thread is mapped to an interior edge. In line 4-9, the edge number is used to obtain the IDs of the two neighboring elements as well as their local edge numbers (1-3), with the interior side denoted with and exterior . Line integration on the edge is done by looping through the edge quadrature points. First, at each point we compute the state variables (Line 14-22), use them to compute the numerical flux. Then, we add this flux contributions to both sides of the edge, indexed using the local edge number of each element (line 34-40). The subsequent elemental flux summation is shown in Algorithm 3, where each thread is mapped to an element and sums the contributions from its 3 edges.
5.2.3 Wind kernel loops
Wind and pressure calculations are done in a series of small loops, each operating on the mesh nodes independently. In the OpenACC version, we simply parallelize each loop with !$acc parallel loop. Similarly in the CUDA version, we can use the Kernel Loop Directive !$cuf kernel do <<< *, * >>> to parallelize each loop. The reason we did not also use this in the integration kernels is that it is much more limited. For example (to our knowledge), it does not allow privatizing automatic arrays as in the OpenACC case. Using the full CUDA API also allows for future optimizations, for example shared memory manipulation, in a straightforward way.
In the CUDA version, we also declare as constant several fixed-sized arrays used in wind interpolation, which results in significant speedup.
5.2.4 Data transfers for I/O
At every fixed number of time steps specified by the user, the program writes the state variables to ascii files. Thus data will have to be copied to the CPU. While it is possible to rely on managed or unified memory system to initiate the transfers, they can incur large overheads. In the CUDA version, as we traverse through the arrays to write the values, data is fetched page-by-page when page faults are triggered. In the OpenACC code, using unified memory results in the same problem, as there are enough access counts to trigger page faults. To avoid this, in both versions we use cudaMemPrefetchAsync to transfer data at once, as shown in Listing 3.
@@ 5.2.5 Memory usage of automatic arrays
In the OpenACC code, we sometimes observe runtime memory error when parallelizing loop with many automatic (private) arrays per thread. From our observation, this seems to be due to the allocation on these arrays on the GPU stack, a different behavior from using the CUDA API. We can avoid this by limiting the number of gangs to a fixed number, e.g.
@@
as opposed to allowing an automatic scheduling of the gangs.
@@ 5.3 Multiple GPUs
For the OpenACC version, we also add support for a multi-GPU setup using CUDA-aware MPI which can be built on top of existing MPI routines. Namely, we modify the packing (gather) and unpacking (scatter) of ghost data to be computed the GPU. This removes data movement between host and device, and takes advantage of the higher memory bandwidth on the GPU.
6 Experimental Results
6.1 Test cases
To observe performance improvements, we run the solver on three realistic flooding scenarios due to Hurricane Harvey (2017): one on the Neches River and two on the whole Gulf of Mexico.
Flooding in Neches River (neches)
The first scenario simulates flooding in the Lower Neches River watershed during Hurricane Harvey. This setup has been used in past compound flooding studies [Wichitrnithed et al. (2024), Loveland et al. (2021)]. The mesh is comprise of 122,839 triangular elements and 62,075 nodes and is plotted in Figure 4. The upstream boundary is implemented using time-varying normal flux, and the downstream boundary a time-varying water elevation. Together they model the impact of river discharge in a compound flooding scenario.
Hurricane Harvey on a small mesh (harvey1)
@@ 6.1 Test cases
This larger mesh is one of the cases in the ADCIRC testsuite and contains 575,512 elements and around 300,000 nodes. We impose a tidal (periodic) elevation boundary on the right side and use Harvey Best Track data from the National Hurricane Center HURDAT2 database.
Hurricane Harvey on a large mesh (harvey2)
This large setup uses the same data for Harvey as in harvey1, but the mesh is significantly more resolved (Figure 5). The mesh has been validated in past hindcasting studies and contains 6,675,517 elements and 3,352,598 nodes with a maximum resolution of 20 meters.
6.2 Results
We measure the wall clock time of each test case on each platform. In each case, we compare single-GPU performance to the baseline CPU MPI implementation using 144 cores, the maximum number for a single CPU node on GH200. To normalize the measurements across the different meshes and timestep size, we convert wall clock time to million elements per second (MEPS) as the metric for performance, calculated as
These values along with the speedups are listed in Table 5 and plotted in Figure 6. Both CUDA and OpenACC versions show similar performance and speedup over CPU code. We observe less speedup in the neches case (2.3-2.5) compared to the two Harvey cases (6.2-7.3) due to the absence of wind forcing (discussed below). Both Harvey cases show similar performances for all versions including CPU, demonstrating a good scalability on the mesh size. Minor differences can also arise from the mesh structure as well as runtime physics parameters.
| Case | No. of elements | 144 CPUs | CUDA | OpenACC |
| Neches river | 122,839 | 91.6 | 232 (2.5) | 210 (2.3) |
| Harvey 1 | 575,512 | 42 | 276 (6.6) | 261 (6.2) |
| Harvey 2 | 6,675,517 | 45.9 | 337 (7.3) | 320 (6.9) |
Next, we break down the runtime contributions from different subroutines, shown in Figure 7 for each dataset. Note that we only include the most expensive subroutines. The meaning of the various subroutine names are described in Table 6. Again, we do not observe significant differences between CUDA and OpenACC, and the most expensive subroutine is the interior edge integration, followed by wind forcing (in the harvey cases), and volume integration.
However, wind forcing takes the most time in the CPU code (60% and 56% for harvey1 and harvey2, respectively) which explains why we see greater speedups in the harvey cases over neches. Slope limiting is particularly expensive in the CPU version largely due to the communication ghost nodal values between neighboring elements across MPI domains.
| Name | Subroutine |
internal_edge |
Interior flux integration + flux gathering |
rhs_dg |
Volume integration |
rk_update |
Runge-kutta update loop for the next stage |
slopelim |
Slope limiter |
wetdry |
Wetting and drying scheme |
met_forcing |
Computation of wind and pressure |
Finally, strong scaling is evaluated for harvey2 (Figure 8) for 1-8 GPUs.
7 Discussion
This section discusses the performance results reported and how well GPU resources are being utilized. We focus on the most time-consuming kernels and discuss the harvey1 case in place of harvey2, as they are similar. Since there are no significant performance differences between CUDA and OpenACC, we only show the analysis for the CUDA version.
An important performance metric we report is the arithmetic intensity , defined as the ratio of the number of floating point operations (FLOPs) to the number of bytes transferred to/from the level in the memory hierarchy:
where can refer to the High Bandwidth (global) Memory (HBM), the L2 cache, and the L1 cache. This number, along with the performance in FLOPs per second, are obtained empirically from the NVIDIA Nsight Compute profiler. It is used in the Hierarchical Roofline Model [Yang et al. (2020)] to obtain performance bounds based on the relation:

### v2

# Wichitrnithed, September 21, 2025 GPU-acceleration of the Discontinuous Galerkin Shallow Water Equations Model (DG-SWEM) with OpenACC

@@ Abstract
This paper presents a porting of DG-SWEM, a discontinuous Galerkin solver for hurricane storm surge, to NVIDIA GPUs. Time-explicit discontinuous Galerkin methods contain a large number of degrees of freedom but have been shown to exhibit a large amount of data parallelism due to the loose coupling between elements, and thus are naturally mapped to the GPU architecture. A previous framework in porting DG-SWEM to GPUs required converting subroutines from Fortran to C++ to be used with CUDA C++. By using OpenACC and Unified Memory, we simplify the porting process and maintain a single codebase for both CPU and GPU versions. We test the code using a large Hurricane Harvey scenario on NVIDIA’s Grace Hopper chip, and compare the GPU code’s performance on multiple H200 nodes to the CPU version on the same amount of Grace CPU nodes.
@@ 1 Introduction
A primary disadvantage of the DG method which has confined DG-SWEM to mainly research and not operational forecasting like ADCIRC is the increased computational cost associated with the computations required from the higher number of global degrees of freedom compared to the continuous formulation. At the same time, the localized nature of computing the solution for each element independently generates a high degree of data parallelism, and several GPU implementations of the DG method have been using different approaches. Early implementations using NVIDIA’s CUDA include [Klöckner et al. (2009), Siebenborn et al. (2012), Fuhry (2013), Brodtkorb et al. (2012)]. More recent implementations using different application programming interfaces (API) can be seen in [Abdi et al. (2019), Kirby and Mavriplis (2020)] which use OCCA, and [Xia et al. (2014), Dai et al. (2022)] which use OpenACC. Lastly, we mention the recent open-source FEniCSx SWE solverby Dawson et al. in [Dawson et al. (2024)] and GPU accelerated by Pachev in [Pachev (2024)], where implicit schemes are considered.
In Section 2, we introduce the governing model, i.e., the SWE and our DG discretization. Next, in Sections 3 and 4, we discuss the related work and the hardware setup we utilize. Section 5 details our implementation. In Section 6, we discuss multi-node performance results when running the code on a large Hurricane dataset. Finally, in Section 7 we draw conclusions and discuss potential future research directions.
@@ 3 GPU Programming approach
As an alternative to purely distributed parallelism, graphics processing units (GPUs) can be viewed as a multithreaded vector processor [Hennessy and Patterson (2018)], where each unit executes the same instructions on multiple data (SIMD), but individual lanes can take different paths. To take advantage of a large number of such processors, the algorithm must exhibit a high degree of data parallelism, where multiple elements can be processed independently at the same time. This is particularly applicable to explicit DG methods due to the inherently local computation of the elements.
In the past, as GPUs have originally been designed for graphics programming, the available application programming interface (API) such as DirectX and OpenGL imposed limitations of using GPUs for general-purpose computing (GPGPU) [Kirk and Hwu (2010)]. In 2007, NVIDIA developed CUDA [NVIDIA (2024)] which provides a more generic API to GPU instructions. This significantly extended the capabilities of NVIDIA GPUs to scientific and engineering codes: the programmer gained control of individual GPU threads, synchronization, and shared memory. Using the NVIDIA HPC compiler stack (a successor of PGI compilers), CUDA is supported in C, C++, and Fortran [NVIDIA (2020)]; CUDA code can be written in the same source files and compiled with appropriate flags. While the low-level nature of CUDA allows more opportunities for control and optimization by the programmer, it typically results in having to maintain a separate code base from the CPU code. In the past, a framework has been created to generate GPU code from DG-SWEM through CUDA C++ [DuChene et al. (2011)]. The workflow consisted of scripts that transform Fortran subroutines to C++ and handle variable declaration and allocation. The work noted that other alternatives at the time, such as using the PGI accelerator or CUDA-Fortran, were lacking in features or accessibility.
A more modern approach to GPU programming is to use OpenACC [OpenACC-Standard.org (2021)] or OpenMP [Openmp.org (2021)]
to write directives instead of manually managing individual GPU threads. This approach allows the programmer to insert high-level statements that generate GPU instructions and appear as comments in the source code (similar to OpenMP loop constructs, for instance). This can sometimes result in a single code base — only the compilation flags such as -acc need to be turned on to enable the GPU code generation. Of course, the programmer is still required to sufficiently expose the parallelism in the code, which often results in modifications of the CPU code as well. We choose to use OpenACC in this work as it currently has better support on the NVHPC compilers.
In this approach, we aim to mimimize integration and maintenance efforts, as well as to take advantage of the new NVIDIA Grace-Hopper architecture [Singer (2024)] which provides a more efficient CPU-GPU data transfer.
@@ 4 Experimental setup
The original CPU code was compiled using the Intel oneAPI compiler stack. This work uses the NVIDIA SDK compiler stack 24.7 (nvfortran, nvc, nvc++) for both CPU and GPU versions. Compilation flags for the CPU and GPU codes are listed in Table 3 .
@@
| GPU | mpif90 -r4 -O3 -cuda -acc -Mextend -Mpreprocess -gpu=unified |
@@ 5 Implementation
5.1 Unified Memory
A significant part of porting code to GPUs is managing data movement. As a CPU (host) and a GPU (device) have separate physical memories, data must be transferred back and forth to perform the respective calculations. Traditionally, even if some arrays are only used in device calculations, they still need to be initialized or read from files through the host, and we would have to declare device copies of those arrays. For example, in CUDA Fortran, an existing array in the code
@@ 5.1 Unified Memory
To avoid having to do this for each array in the code, we exploit
Unified Memory which was introduced in CUDA 6.0. This allows GPU kernels to access data on the host without needing to explicitly copy the data to device. Data migration is automatically performed once a page fault is triggered on the device. In the past, this requires a special allocation, e.g. cudaMallocManaged in CUDA C.
For newer systems that support Heterogenous Memory Management (HMM), all device data declarations can be completely eliminated using the compilation flag -gpu=unified. This significantly simplifies programming effort and enhances maintainability. In both approaches, a synchronization statement like !$acc wait is needed for correctness before the CPU can access the computed data, since GPU kernels immediately return control to the host.
In Table 4 we summarize the important data arrays, some of which are referred to in the kernel descriptions in the next section.
| Array | Meaning |
| ZE[k,e,irk] | mode of elevation solution at element on RK stage irk |
| QX[k,e,irk] | mode of x-velocity solution at element on RK stage irk |
| QY[k,e,irk] | mode of y-velocity solution at element on RK stage irk |
| RHS_ZE[k,e,irk] | Contribution to the mode of at element on RK stage irk |
| RHS_QX[k,e,irk] | Contribution to the mode of at element on RK stage irk |
| RHS_QY[k,e,irk] | Contribution to the mode of at element on RK stage irk |
| PHI_AREA[k,i] | basis evaluated at area quadrature point i |
| PHI_EDGE[k,i] | basis evaluated at edge quadrature point i |
| XFAC[k,i,e] | Weight for the basis of element in x-direction at area quadrature point i |
| YFAC[k,i,e] | Weight for the basis of element in x-direction at area quadrature point i |
| SRFAC[k,i,e] | Weight for the basis of the source term on element at area quadrature point i |
| EDGEFAC[k,i,l] | Weight for the basis of the source term on element at area quadrature point i |
| EDGE_ze[k,l,e] | Numerical flux contribution to the basis of at local edge of element |
| EDGE_qx[k,l,e] | Numerical flux contribution to the basis of at local edge of element |
| EDGE_qy[k,l,e] | Numerical flux contribution to the basis of at local edge of element |
@@ 5.2 GPU kernels
This section outlines the main subroutines and how their parallelization. For the integration kernels, we follow the one-thread-per-element and one-thread-per-edge approach as in [Fuhry (2013)].
@@ 5.2.1 Area integration
The area integration routine (Algorithm 1) is naturally data-parallel, as computations are done locally to each element . At a given Runge-Kutta stage irk, it computes the terms
from (3) and accumulates them into the RHS vectors. Recall from Eq 2 that the flux terms (explicitly written in terms of the state variables) are
@@
There are technically three levels of loop parallelism: looping over the elements, looping over the quadrature points, and looping over the degrees of freedom. Here, the element-level loop is accelerated through the gang-vector parallelism over the number of elements (line 2), and the inner loops over quadrature points and degrees of freedom are enforced to be serial using !$acc seq. This needs to be explicitly inserted, as by default the inner loops will also be parallelized, and in this case the trip counts is too short to be profitable. We also add the private(...) clause to keep any local arrays private to each thread.
Importantly, the async clause is added to eliminate kernel launch overhead, and the default(present) the overhead of checking if data is already on the device. This same structure is applied for the edge integration kernel. We have observed that using these clauses speeds up the program by almost 1.5.
@@ 5.2.2 Edge integration
In Algorithm 2, each GPU thread is mapped to an interior edge, again through gang-vector parallelism (line 2). In line 4-9, the edge number is used to obtain the IDs of the two neighboring elements as well as their local edge numbers (1-3), with the interior side denoted with and exterior . Line integration on the edge is done by looping through the edge quadrature points. First, at each point we compute the state variables (Line 14-23), use them to compute the numerical flux. Then, we add this flux contributions to both sides of the edge, indexed using the local edge number of each element (line 35-41). The subsequent elemental flux summation is shown in Algorithm 3, where each thread is mapped to an element and sums the contributions from its 3 edges. Note that for brevity, we have opted for the atomic approach for other less common boundary integration routines.
5.2.3 Other kernels
Other kernels like slope limiting (SLOPELIM) and wet/dry calculations (WETDRY) are similarly parallelized at the element-level loop, with small inner loops serialized. Wind forcing and pressure calculations (WIND) are done in a series of loops, each operating on the mesh nodes independently. We simply parallelize each loop with !$acc parallel loop.
5.2.4 Data transfers between host and device
At every fixed number of time steps specified by the user, the program writes the state variables to ASCII files. Thus data will have to be copied to the CPU. While it is possible to completely rely on HMM to initiate the transfers, they can incur large overhead as there are enough access counts to trigger page migration. To avoid this, we use cudaMemPrefetchAsync to transfer data at once, as shown in Listing 1.
@@ 5.2.5 Memory usage of automatic arrays
We sometimes observe runtime memory error when parallelizing loop with many automatic (private) arrays per thread. From our observation, this seems to be due to the allocation on these arrays on the GPU stack, a different behavior from using the CUDA API. We can avoid this by limiting the number of gangs to a fixed number, e.g.
@@
as opposed to allowing an automatic scheduling of gangs.
@@ 5.3 Multiple GPUs
To exchange data among different domains when using multiple processors, DG-SWEM uses MPI persistent communication. At the start of the simulation, calls to MPI_SEND_INIT() and MPI_RECV_INIT() are made to specify the location of data to exchange to each neighbor. Then, as shown in Listing 2, whenever we need to exchange data, we call MPI_STARTALL() to initiate the communication in a non-blocking manner. At the end, we wait until all the messages have arrived. The packing of the elements to send is done through the subroutine packElems(), and the unpacking similarly through unpackElems().
The existing MPI routines need no further modification in the GPU code, as they accept unified memory addresses. However, we observe data being transferred to host and back to device during MPI communication. This appears to be a limitation of using unified memory with CUDA-aware MPI which prevents direct GPU-to-GPU data transfer [Chenbr (2025)]. The packing and unpacking have been parallelized to take advantage of the higher memory bandwidth of the GPU, as shown in Listing 3.
6 Results
6.1 Test Case
@@
To measure performance improvements, we run the solver on a dataset for Hurricane Harvey (2017) which is representative of a typical simulation scenario. The mesh is shown in Figure 2 which contains 6,675,517 elements and 3,352,598 nodes with a maximum resolution of 20 meters. We impose a tidal (periodic) elevation boundary on the right side and use Harvey Best Track data from the National Hurricane Center HURDAT2 database. The timestep size is set to be s, and computation is done with single-precision.
Figure 3 shows the performance comparison between the GPU implementation and the CPU implementation from 1 to 8 nodes, where we run the simulation for 28,800 time steps. As described in Section 4, each GPU node consists of a single NVIDIA H200, while each CPU node contains 144 NVIDIA Grace cores. Here, ”performance” is measured by the rate of work done, defined to be the number of degrees of freedom computed per second. In a node-by-node comparison, we observe a speedup of GPU code over CPU of around to . Figure 4 further shows the breakdown of time spent in each kernel. In this case, we only consider the total time spent in the timestepping loop (i.e. exclude I/O or startup time) and run the simulation for 2,880 time steps. Here, we see that parallel communication time in the GPU case increases with more GPUs to around 6 s at 8 nodes, resulting in a scaling plateau. For the same amount of nodes in the CPU case case, the communication time appears to stabilize to around 18 s.
6.2 Performance Evaluation
To further evaluate performance of the time-consuming kernels, we first calculate the theoretical arithmetic intensity at the DRAM level. This is the ratio of floating point operations (FLOPs) to total memory traffic in bytes from/to DRAM:
We count each arithmetic operation like addition and division counts as one FLOP, including special operations like square root and power. In estimating from the source code (by manually counting), we also assume that everything is loaded once from memory for simplicity, which is generally not valid for large enough datasets and when accesses are not fully coalesced. Thus we expect the estimated values to be higher than the true values. The estimated arithmetic intensities for the integration kernels are shown in Table 5, along with their experimental values obtained from the NVIDIA Nsight Compute profiler. We focus on these two kernels as they are representative of the DG method.
| Estimated | Experimental | |||
| Kernel | GFLOPs | GBytes | ||
| EDGE | 7.95 | 1.6 | 4.97 | 2.7 |
| AREA | 4.09 | 2.2 | 1.85 | 1.81 |
| GATHER | 0.24 | 1.92 | 0.125 | 0.09 |
It can be seen that the arithmetic intensity of AREA is very close to the estimated value. This is expected, as most data are not shared between different elements, and are loaded in a (mostly) coalesced manner. A similar reasoning applies for GATHER. On the other hand, EDGE has a markedly lower intensity than estimated, a little over half. This can be attributed to the different memory access patterns. While area integration loops over elemental arrays sequentially, edge integration requires each edge to fetch data from neighboring elements which may not lie close together in memory. This poor spatial locality is a well-known issue in using unstructured meshes [Hadade et al. (2020)] and in our case results in uncoalesced accesses and therefore more data traffic.
We further evaluate the performance using the empirical Roofline which prescribes a performance bound through the relation: